Can you please show step by step and also how to do it on HP financial calculator.

Your firm recently paid a dividend of $4 to common stockholders. Dividends are expected to grow at 8% per year for the foreseeable future. The current stock price is $54.

Preferred stock would pay a 12% dividend on a $50 par value. The stocks would sell for par value less flotation costs of $2 per share. Wellington has a marginal tax rate of 34%.
What is the firm’s cost of capital if their capital structure consists of 60% equity and 40% preferred stock?

The answer is 14.60% but idk how to do it
Answer #1

Given that,

For common equity,

last dividend D0 = $4

growth rate g = 8%

current price P0 = $54

So, cost of equity using constant dividend growth model is

Cost of equity Ke = D0*(1+g)/P0 + g = 4*1.08/54 + 0.08 = 0.16 or 16%

For preferred stock,

par value = $50

dividend = 12% on par = 0.12*50 = $6

Flotation cost = $2

So, cost of preferred stock Kp = Dividend/(Par - Flotation cost) = 6/(50-2) = 12.50%

Weight of equity We = 60%

Weight of preferred stock Wp = 40%

So, Cost of capital = We*Ke + Wp*Kp = 0.6*16 + 0.4*12.5 = 14.60%